0.00325 times 10^-8 can also be expressed in mm a
Answer #1

1) 1cm = 10 mm

0.00325 x 10 -8 cm = 3.25 x 10-11 cm = 3.25 x 10-11 x 10 mm = 3.25 x 10-10 mm

Option C

2) Option A

ay = - g where g= 9.8 m/s2

Eaath's gravitational force is constant and is pointed in downward direction. Hence ay is negative.

3) Let P and Q are the end points of vector C

Co ordinates of P and Q are : P = (1,-1) Q = (-3,1)

x and y components of vector C = Q-P = (-3 -1 , 1 +1) = (-4 , 2)

Option : E

4) Total flight time of the projectile (2u/g) depends only on the speed . Hence the flight time for both projectiles is same.

Option : C

5) Option B

The car speeds on the incline after a striaght road

6) The inclined force T upward has a component in Y- direction facing upward .So the net normal force on the body will reduce and is less than mg

Option C n<mg

7) Each string exerts a force of W N to suspend the hanging mass to which it is attached. This means the tension in the string is W N. So the scale will read W N.

Option : A

8) By definition , In an INEALSTIC collision, momentum is conserved but kinetic enrgy is not conserved. The difference in kinetic energy is converted to thermal energy during collision.

Option : C

9) Option B

Normal force is equal to the weight at anuy instant.

10) option D

Momentum is conserved